Upon graduating from college, you make an annual salary of $47,892. You set a goal to double it in the future. If your salary increases at an average annual rate of 6.57 percent, how long will it take to reach your goal?

Round the answer to two decimal places.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

(2*47,892)=47,892*(1.0657)^n

2=(1.0657)^n

Taking log on both sides;

log 2=n*log (1.0657)

n=log 2/log (1.0657)

=10.89 years(Approx).